# Re-examining learning linear functions in context

### v1

@@ 4.4 Predictions for models with only attention layers or with only MLP
To understand better which components in the transformer architecture are responsible for ICL, we tested various components. We found that attention layers (AL) were the important components for ICL but ICL only worked reasonably well when the model had 2 AL (see also figure 3). Beyond 2 AL what mattered most was the number of attention heads (whether they are summed over all layers or counted within a layer). A single AL model had only a very limited ICL generalization capability beyond testing on $D^{t}_{F}=N(0,1)$, but it did better than a 12 layer MLP, which showed no ICL capability, probably because the method of training on the predict next token format is not suitable for models without attention heads. The details of various AL models are found in Figure 5 and Table 3.
@@
4.5 Ordering and restricting the size of prompts can improve performance
Our experiments with distributions also showed that a model performance improves when the sequence of prompts for the $x_{i}$ follows or are “sorted” to follow the natural order on $\mathbb{R}$, especially for bigger models. Error rates were comparable to the original models without sorting for small test values of $\sigma$ with $D^{t}_{F}=N(0,\sigma)$ and substantially lowered error rate, by up to a third depending on the training distribution, from the unsorted models. The details concerning the sorted models are in Table 1.
While at least 2 points are needed to find a linear function, we noticed that model performance degrades when the size of the prompt during inference is greater than the maximal size of prompts the model saw during training, as the rightmost plot in Figure 6 shows (Appendix C). This held over all models and training distributions. This means that the model takes into account the whole sequence in its calculations, not just the last two or three data points. Had the model only looked at a small fixed subsequence, large size prompts in inference would not have affected model behavior.
4.6 Taking stock: what are the models learning
To ICL  ${\cal L}$, we expected a transformer model given $(x_{1},f(x_{1}),...,x_{n}),?)$ to perform a linear regression. The hypothesis and theoretical constructions of Akyürek et al. (2022); Von Oswald et al. (2023) shows that transformer models can perform linear regression with 1 step of gradient descent. If that were the case, the models should generalize without difficulty. But this is clearly not what we observed. Error rates depend on training distributions $D_{F},D_{I}$ and on the distance of the target function’s values from the majority of the data points in the model’s training. We’ve also demonstrated model sensitivity to the entire sequence of ICL prompts. This shows that the models did not learn to use linear regression to solve this task. All our models failed to learn the concept of a strictly monotone increasing or decreasing linear function in ${\cal L}$ over arbitrarily large or at least many large intervals of $\mathbb{R}$ [^3] This makes sense in terms of Asher et al. (2023)’s characterization of learnability. The concept of a strictly monotone increasing or decreasing linear function describes a $\Pi^{0}_{1}$ set in the Borel hierarchy which Asher et al. (2023) show is not learnable using ordinary LLM assumptions.
@@ 4.6 Taking stock: what are the models learning
Our error analysis has shown us the existence of boundary values, values for which models do well on the interval defined by the boundary values but degrade outside of them. These boundary values fluctuate depending on model training distributions. This is further evidence that the model’s method for solving the task does not involve a real calculation of a linear function but an adjustment of values that the model has seen.
5 How might the models be learning?
Olsson et al. (2022) argue that a copying and comparison mechanism (induction head) is at the heart of ICL. They show that induction heads do not exist for 1 layer attention-only models but do for attention-only models with two or more layers. They empirically establish a strong connection between the formation of induction heads in models with greater than 2 attention layers and the model’s ability to ICL. Our experiments with attention-only models showed that multiple attention layers were needed for ICL. Attention-only models with induction heads could ICL linear functions reasonably well, at least in when $D_{F}=D^{t}_{F}$; and in fact the large 2 attention only layer model with 32 AH was more robust than the full transformer model with 1 (attention and MLP layer) and 1 or 2 AH (See Table 2 Appendix B). Olsson et al. (2022) also note that induction head behavior is possible with 1 attention + MLP layer.
@@ 5 How might the models be learning?
Our induction head hypothesis is that the induction heads predict a value for $f(x_{n})$ given a prompt sequence $\vec{x}=(x_{1,1},x_{1,2}(=f(x_{1})),x_{2,1},x_{2,2},...x_{n,1},?)$ by using a projection from similar sequences or subsequences in the training, $\vec{y}=(y_{1,1},y_{2,2}...y_{n,1},y_{n,2})$, with $x_{i,1}$ close to $y_{i,1}$ for some $j$ and $x_{i,2}$ close to $y_{j,2}$. Given the effects of prompt length on performance, we know that the whole sequence matters with $p_{2}\leq p_{1}$ for optimal predictions. This is evidence for a pointwise comparison like we are proposing (which is more complicated and potentially more accurate than simply averaging the $y_{n,2}$ of the three closest $y_{n,1}$ neighbors of $x_{n,1}$).
Olsson et al. (2022) report that larger models’ induction heads can exploit sequences $\vec{y}$ that are “more dissimilar” to each other than smaller models can. The fact that the larger models respond well to prompts ordered according to the natural ordering on $\mathbb{R}$ suggests that larger models with more attention heads can exploit comparing sequences that converge or diverge from the target sequence $\vec{x}$ as the prompts $x_{i,1}$ near $x_{n,1}$ increase or decrease. And they can compare in different ways. For smaller models that did not improve performance with respect to naturally ordered prompts, the hypothesis suggests they are restricted to simpler operations on the sequences they have seen. This and our observations about boundary values provide further empirical support for the induction head hypothesis.
Given boundary values, $-B,B$, all or the vast majority of the sequences the model has seen have values $z_{i}$ with $-B<z_{i}<B$. If the target sequence $\vec{x}$ has maximum values $-B<x_{i}<B$, i.e. $-B<Maxval_{x_{i}}\vec{x}<B$, then chances are high that the model will find a weighted set of sequences $Y$ close to the test sequence $\vec{x}$ and compute bounds for $x_{n,2}=f(x_{n}))$. Call the sequence $\vec{y}$ generated by a function $g=a_{g}x+b_{g}$ a g-sequence. We assume the standard measure over sequences. Given a g-sequence $\vec{y}$ closest to the f-sequence $\vec{x}$ generated by target function $f=a_{f}x+b_{f}$ is such that $a_{g}=a_{f}$, then the model will be able to approximate $x_{n,2}$ by averaging the distances between $y_{i,2}$ and $x_{j,2}$ for the closest $y_{i,1}$ to $x_{j,1}$ for all $x_{j,1}$. Now suppose that the closest g-sequences $Y$ are not all such that $a_{f}=a_{g}$. The model must now construct a function $\mathfrak{h}(Y_{\vec{x}},\vec{x})$ that computes a distance $d$ between the values it has seen in $Y_{\vec{x}}$ and the targets $\vec{x}$ for some optimized set $Y_{\vec{x}}$ of sequences close to $\vec{x}$. If $\mathfrak{h}(Y_{\vec{x}},\vec{x})(x_{k,1})=z_{k,2}$ is the k-th member of $\mathfrak{h}(Y_{\vec{x}},\vec{x})$, we optimize $\mathfrak{h}$ such that $|z_{k,2}-x_{k,2}|$ is minimized for all $k$. The model then averages these distances to yield an ”average” $\mathfrak{h}(Y_{\vec{x}},\vec{x})$ to compute $z_{2,n}=\hat{f}(x_{1,n})$. The larger the set very close $\vec{y}\in Y_{\vec{x}}$, the better the projection and hence the prediction.
For prompts outside the boundary values $-B,B$, the closest $\vec{y}$ are those with values near the boundary ($y_{n,2}\approx B(-B)$). So the model will predict $x_{n,2}\approx B(-B)$, if it uses this method of projection. So in sum, we think the model estimates a set $Y$ of closest sequences to the target $\vec{x}$ and computes:
$\hat{f}(x_{n})=x_{n,2}=\frac{1}{n}\sum^{n}_{i=1}\mathfrak{h}(Y_{\vec{x},x_{i}})(x_{n,1}),\ \mbox{\em for}\ -B<Maxval_{x_{i}}\vec{x}<B$
[…]
$\mbox{\em Otherwise}\hat{f}(x_{n})\ \mbox{\em takes a random value}\in[-B,B],\ \alpha_{\cal L}>0\ \mbox{\em a characterstic model value}$
The observations we have made are clearly compatible with this hypothesis (See Table 10 Appendix F) and the weighted averages are calculable in a 2 layer Attention only model with suitable heads. This also explains why training the model with smaller sequences in a kind of ”curriculum learning” may not be helpful. We trained the same 12L8AH model once with curriculum learning which supposedly helps the model perform better for different types of size prompts and generalize. and the same without curriculum and found that the model without curriculum performs better. This is consistent with our hypothesis, since the model without curriculum looks at batch sizes of up to 41 points in each step, and can therefore see more sequences than the model with curriculum, which looks at batch sizes of up to 41 points. The induction head hypothesis is less precise then linear regression but can approximate it given an appropriate set $Y$. Our induction head hypothesis predicts that model performance will be sensitive to a choice of training distribution for $D_{F},D_{I}$ as well as a choice of test distributions and of course to the presence of boundary values.
@@ 6 Conclusion
Appendix A: Training details
@@
Appendix B: Table of error progression for models trained on $N(0,1)$ distributions tested on $N(0,\sigma)$
@@
Appendix C: Failure to generalize to longer prompt sequences
Appendix D: Plots for boundary values with $U(-5,5)$
@@
Appendix E: Example of boundary values for attention only models
@@
Appendix F: The model searches for a sequence close to the input sequence.
Figure 10: Plots model 12L8AH, trained on $D_{I}=D_{F}=N(0,1)$ for $f(x)=x$ for high values of $x$ and $f(x)=10x$ for low values of $x$

### v2

@@ 4.4 Predictions for models with only attention layers or with only MLP
To understand better which components in the transformer architecture are responsible for ICL, we tested various components. We found that attention layers (AL) were the important components for ICL but ICL only worked reasonably well when the model had 2 AL (see also figure 4). Beyond 2 AL what mattered most was the number of attention heads (whether they are summed over all layers or counted within a layer). A single AL model had only a very limited ICL generalization capability beyond testing on $D^{t}_{F}=N(0,1)$, but it did better than a 12 layer MLP, which showed no ICL capability.Attention-only models could ICL linear functions reasonably well, at least in when $D_{F}=D^{t}_{F}$; the large 2 attention only layer model with 32 AH was more robust than the full transformer model with 1 (AL and MLP layer) and 1 or 2 AH (See Table 2 Appendix B). Tables 3 4 in Appendix and Figure 3 give details of various AL models on normal and uniform distributions. It may be interesting to look at the results for attention layers when they are provably causally effective, as with efficient attention Naim & Asher (2024).
@@
4.5 Ordering prompts and restricting their size
Model performance improves when the sequence of prompts for the $x_{i}$ are sorted to follow the natural order on $\mathbb{R}$, especially for bigger models. Error rates compare to error rates without sorting for small values of $\sigma$ with $D^{t}_{F}=N(0,\sigma)$ and are lower by up to a third on other test values, depending on the training distribution (see Table 1).
While at least 2 points are needed to find a linear function, all model performance regardless of training distribution degrades when the size of the prompt during inference is greater than the maximal size of prompts seen in training, as the rightmost plot in Figure 9 shows (Appendix E). Further models did better with the distributions that were exactly the size (41 data points) of those in their training We tested a 12L8AH model with with smaller sequences in a kind of ”curriculum learning” and without curriculum; we found that the model without curriculum training performed better. All this implies that a model takes into account the whole sequence in its calculations, not just the last two or three data points. Had the model only looked at a small fixed subsequence, larger sized prompts in inference would not have affected model performance and curriculum learning should have improved it.
5 What and how are the models learning?
The hypotheses and theoretical constructions of Akyürek et al. (2022); Von Oswald et al. (2023) led us to expect that a transformer model given $(x_{1},f(x_{1}),...,x_{n}),?)$ would perform a linear regression to ICL a linear function. In this case, the models should generalize without difficulty. But this is not what we observed. Error rates depend on the distance of the target function’s values from the majority of the data points in the model’s training. Models are also sensitive to the entire sequence of ICL prompts, not just the minimal number needed to compute a linear function. Error analysis showed the existence of boundary values $-B,B$; models do well on the interval $[-B,B]$ degrade outside of them. These boundary values fluctuate depending on model training distributions and size. All this is strong evidence that models did not learn to use linear regression to solve this task and failed to learn the concept of a strictly monotone increasing or decreasing linear function in ${\cal L}$ over arbitrarily large or at least many large intervals of $\mathbb{R}$ [^2] This makes sense in terms of Asher et al. (2023)’s characterization of learnability. The concept of a strictly monotone increasing or decreasing linear function describes a $\Pi^{0}_{1}$ set in the Borel hierarchy which Asher et al. (2023) show is not learnable using ordinary LLM assumptions.
@@ 5 What and how are the models learning?
Rather than computing a linear function in this task, the models estimate continuations of sequences based on sequences they have seen. This is in line with Olsson et al. (2022)’s finding that a copying and comparison mechanism (induction head) is at the heart of ICL. They show that induction heads only exist for attention-only models with two or more layers and that larger models’ induction heads can exploit sequences that are “more dissimilar” to each other than smaller models can.
@@
Our induction head hypothesis is that a model predicts the value for a given a prompt sequence by using a projection from similar sequences or subsequences in the training. Our observations about boundary values provide further empirical support for this assumption. In addition, The effects of prompt length on performance imply that the whole sequence matters with $p_{2}\leq p_{1}$ for optimal predictions.
@@ 6 Conclusion
Appendix A Training details
@@
Appendix B Error progression for models trained on $N(0,1)$ distributions tested on $N(0,\sigma)$
When $D_{I}=D_{F}=N(0,\sigma)$ there is for $x\in N(0,\sigma)$ an over 85% chance of $f(x)\in[-4\sigma^{2}-2\sigma,4\sigma^{2}+2\sigma]$ and a 95% chance $f(x)\in[-2\sigma,2\sigma]$. So a model with $\sigma=1$ $D_{F}=D_{I}=N(0,1)$ has seen sequences of values for $f$ with $f(x)\in[-2,2]$ more than 95% of the time.
@@ Appendix B Error progression for models trained on $N(0,1)$ distributions tested on $N(0,\sigma)$
Appendix C Plots for boundary values with $N(0,1)$ and $U(-5,5)$
@@
As shown in the left plot in Figure 5, $\hat{f}^{+}(v)\approx 30$ for values $v$ for which the ground truth target function $f$ is such that $30\leq f(v)$, and the model predicts an approximally constant function $\hat{f}^{-}(v)\approx-30$ for values $v$ on which $f(v)\leq-30$. Given a training on $U(-5,5)$ we can calculate 30 and -30, with $30=5*5+5$ and $-30=-5*5-5$, to be the boundary values for the models there.
@@
Figure 7: Plots for $f(x)=10x$ by a 12L8ah model and by a 6L4ah model.
Appendix D Example of boundary values for attention only models
@@
Appendix E Failure to generalize to longer prompt sequences